write the equation in the form y=mx+b
(8,-2) and (0,5)
(-8,0) and (5,-2)

Answer by Boreal(15235)

first one slope is change in y /change in x or 7/-8
point slope formula y-y1=m(x-x1), m slope, (x1, y1) point
y-5=(-7/8)x
y=-(7/8)x+5
----------------------
second slope is -2/13
y=-(2/13)(x+8)=(-2/13)x-16/13. One may use either point; a value of 0 makes it easier.